# EFEAR-4D: Ego-Velocity Filtering for Efficient and Accurate 4D radar Odometry

### v1

@@ IV-A2 Comparative evaluation on NTU Dataset
We compare our methods with 4DRadarSLAM and livox_horizon_loam on NTU dataset [28]. Compared with MSC dataset, the number of sequences on NTU dataset is fewer and the scenes are less challenging since there are no dynamic objects also with relatively low speed. As the dataset only provides Livox horizon LiDAR instead of high-end scanning LiDAR, we use livox_horizon_loam as LiDAR odometry baseline. We evaluate them with the rotation part and the translation part of RPE, respectively. The results are shown in Tab. II. In loop 1 and loop 2, which are collected by car with middle speed, the drift of our EFEAR-4D in the Z-axis is much lower than that of 4DRadarSLAM due to the max-intensity point selection. In cp, a low-speed sequence where travel around buildings, our EFEAR-4D performs well as it can extract stable features such as walls and ground. However, in Garden and Nyl, where the scenes are grassland or bush, EFEAR-4D performs slightly worse as the features extracted are not distinguishable enough.
@@ IV-A3 Ego-velocity estimation evaluation
Figure 9: Ego velocity error comparison between REVE and our method on Dynamic 2 sequence.
@@
Figure 10: Scenes of different dynamic and static sequences.
@@
To further evaluate the performance of our ego velocity estimation, we extract 4 sequences on the MSC dataset, where two are in a dynamic environment, and the other two are in a static environment. Scenes are illustrated in Fig. 10. We compare our method with REVE [17], an open-source algorithm using RANSAC and linear least square method to compute ego velocity. As shown in Tab. III, in dynamic environments, the performance of our method is much better than REVE. The estimation errors of ego velocity in Dynamic 2 sequence are depicted in Fig. 9. Our ego velocity estimation has not only smaller errors but also smaller jitter. This is because, compared to REVE, we propose and use the physical model relating Doppler velocity to the real velocity.
@@
Figure 12: Scenes and satellite imagery collected in our dataset.
@@ IV-A4 Self-collected dataset
The route and scenes are depicted in Fig. 12. All the sequences are collected in University Town, Shenzhen, China. Our mobile robot platform is equipped with an Oculii Eagle 4D radar, a stereo camera, an IMU, a Robosense LiDAR, and a ByNav GNSS navigation system. Ground truth trajectory is obtained from GPS/IMU + RTK. It is clear to see that the sensor height has great impact on the number of radar points, as shown in Fig. 13. At the same GPS location, putting the radar sensor at different heights results in various signal behaviours in the urban environment. Surprisingly, the number of radar points is inconsistent at different places, going up and down from place to place.
Sequence 1 depicts the vicinity of a teaching building at night. Given its surroundings of taller buildings, its characteristics remain stable. Thus, when the radar is positioned at a higher height, it captures a large amount of building-related information, resulting in a higher count of point cloud points. Most point cloud acquired at lower heights pertain to adjacent grass areas within this sequence. Data captured at higher heights not only contains fewer points but also fewer discernible features. Insufficient points within a specific frame lead to higher drift, particularly along the Z-axis, resulting in increased errors as illustrated in Tab IV.
In Sequence 2, the point cloud contains a minimal number of points irrespective of the radar’s height, indicating a degraded radar scene characterized by the absence of discernible features. Error rates are lower when the radar is positioned at a lower height. This is because the scene lacks distinctive features and being closer to the ground enables more effective data extraction, particularly on highways. However, in open environments at higher heights, where point cloud are reflected off trees, features are lacking. Moreover, in frames with fewer than 500 points, feature extraction becomes entirely ineffective. Lowering the radar’s height facilitates better feature extraction, resulting in more accurate results as the point cloud diminishes in size.
The scenario depicted in Sequence 3 entails driving along a roadway. Despite notable differences in the point cloud density between high and low settings, the lower setting still provides adequate number of features for our EFEAR-4D to calculate the pose. In such a case, the disparity between the two settings is not significant.
@@
Some findings in our experiment show that 4D radar may show some weakness in terms of perception for certain scenarios. However, the shortcoming are expected to be addressed with the integration of other complementary sensors. In the future, we will explore sensor fusion with 4D radar and cameras to provide a more comprehensive odometry solution.

### v2

@@ IV-A2 Comparative evaluation on NTU Dataset
We compare our methods with 4DRadarSLAM and livox_horizon_loam on NTU dataset [32]. Compared with MSC dataset, the number of sequences on NTU dataset is fewer and the scenes are less challenging since there are no dynamic objects also with relatively low speed. As the dataset only provides Livox horizon LiDAR instead of high-end scanning LiDAR, we use livox_horizon_loam as LiDAR odometry baseline for reference. We evaluate them with the rotation part and the translation part of RPE, respectively. The results are shown in Tab. II. In the cp sequence, a low-speed sequence involving travel around buildings, our EFEAR-4D performs well due to its ability to extract stable features such as walls and the ground. However, in the Garden and Nyl sequences, which consist of grassland or bushes, EFEAR-4D performs slightly worse as the extracted features are not sufficiently distinguishable. In Loop 1 and Loop 2, collected by car at moderate speeds, our drifts are also slightly lower than those of 4DRadarSLAM.
@@ IV-A3 Ego-velocity estimation evaluation
Figure 9: Ego velocity error comparison between REVE and our method on Dynamic 2 sequence. The figures above and below illustrate the estimation error in the x-direction and y-direction over time, respectively.
@@
Figure 10: Scenes of different dynamic and static sequences. Specifically, the Dynamic 1 and Dynamic 2 sequences are extracted from the URBAN_F0 and URBAN_D0 sequences respectively, and the Static 1 and Static 2 sequences are derived from URBAN_A0 and RURAL_D2 in the MSC dataset.
@@
To further evaluate the performance of our ego velocity estimation, we extract 4 sequences from the MSC dataset, where two are in a dynamic environment, and the other two are in a static environment. Scenes and the specific information about these sequences are illustrated in Fig. 10. We compare our method with REVE [19], an open-source algorithm using RANSAC and linear least square method to compute ego velocity. As shown in Tab. III, in dynamic environments, the performance of our method is much better than REVE. The estimation errors of ego velocity in Dynamic 2 sequence are depicted in Fig. 9. Our ego velocity estimation has not only smaller errors but also smaller jitter. This is because, compared to REVE, we propose and use the physical model in (2) that the $x$-axis and $y$-axis components of Doppler velocities from static objects form a circle to estimate ego velocity, which enhances its consistency.
@@
Figure 12: Scenes and satellite imagery collected in our dataset. The scenes include not only structured environments with dense buildings, but also semi-structured environments with more vegetation and fewer buildings, such as campus main roads. Additionally, the sequences of the same paths with different radar altitudes were collected for further study.
@@ IV-A4 Self-collected dataset
The route and scenes are depicted in Fig. 12. All the sequences are collected in University Town, Shenzhen, China. Our mobile robot platform is equipped with an Oculii Eagle 4D radar, configured with a resolution of 0.86m in range, 0.44 deg in azimuth, 0.175 deg in elevation, and 0.27 m/s in Doppler-velocity. Additionally, it features a stereo camera, an IMU, a Robosense LiDAR, and a ByNav GNSS navigation system. Ground truth trajectory is obtained from GPS/IMU + RTK. It is clear to see that the sensor height has a great impact on the number of radar points, as shown in Fig. 13. At the same GPS location, putting the radar sensor at different heights results in various signal behaviours in the urban environment. Surprisingly, the number of radar points is inconsistent at different places, going up and down from place to place.
Sequence 1 captures the vicinity of a teaching building at night. Surrounded by taller buildings, its characteristics remain stable. When the radar is positioned higher, it captures more building-related information, resulting in a denser point cloud. Conversely, radar at lower heights primarily captures adjacent grass areas. Data collected at higher altitudes contain fewer points and discernible features, leading to higher drift along the $z$-axis and increased errors as shown in Table IV.
In Sequence 2, the point cloud contains few points regardless of radar height, indicating a degraded radar scene with few discernible features. Error rates are lower when the radar is positioned lower, as being closer to the ground allows for more effective data extraction, particularly on highways. In open environments at higher heights, the point cloud reflects off trees, resulting in lack of features. Frames with fewer than 500 points lead to ineffective feature extraction. Lowering the radar height facilitates better feature extraction, resulting in more accurate results.
Sequence 3 involves driving along a roadway. Despite differences in point cloud density between high and low settings, the lower setting still provides enough features for EFEAR-4D to calculate the pose. Thus, the disparity between the two settings is not significant.
@@
Some findings from our experiment indicate that 4D radar exhibits weaknesses in perception in certain scenarios, including instability in point cloud density and weak ground reflection due to its low power and limited FOV. These shortcomings can be mitigated by integrating complementary sensors. Future work will explore sensor fusion with 4D radar and cameras to provide a more comprehensive odometry solution.